# Relaxation time approximation revisited and pole/cut structure in retarded correlators

### v2

@@ Abstract
In this paper, we aim to discuss the following two questions:
  • Q1. Mathematical justification of the model of relaxation time approximation (RTA);
  • Q2. Pole/cut structure in the retarded correlators within the kinetic description.
We find that only the RTA with an energy-independent relaxation time can be justified in the case of hard interactions. Accordingly, we propose an alternative approach to restore the collision invariance lacking in traditional RTA. Furthermore, hydrodynamic poles are the long-lived non-analytical structures in this case. Whereas for soft interactions, commonly encountered in relativistic kinetic theory, the gapless eigenspectrum leads to dominant long-lived branch-cuts. When particles are massive or the perturbations are inhomogeneous, the non-analytical structures become more complex and richer.
@@ I Introduction
Over the past several decades, relativistic hydrodynamics has notably contributed to characterizing the dynamic evolution of Quark-Gluon Plasma (QGP), as observed in heavy-ion collision experiments at the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) [1]. Within the framework of phenomenological modeling for these collisions, the initial temporal parameter, crucial for initiating hydrodynamics, has been meticulously minimized to match experimental observations. This adjustment suggests that the system may rapidly equilibrate. Moreover, the applicability of hydrodynamics has been noted to extend into less expected domains, including small collision systems like nucleus-nucleon and proton-proton collisions, as evidenced by recent studies [2, 3]. What is the dynamic mechanism that triggers rapid equilibration? Why does relativistic hydrodynamics also work well for small systems? Regarding the first question, it has been proposed that hydrodynamics can be viewed as an attractor governing the late-time behavior of systems as they approach equilibrium [4]. At the early stage of evolution, the system tends to flow towards the hydrodynamic attractor, even when far from local equilibrium, which may account for rapid hydrodynamization. Thus, hydrodynamization may have a broader application range than thermalization, contrasting with the conventional view that hydrodynamics is a truncated gradient expansion near local equilibrium. Despite significant advancements, it remains an open question how hydrodynamization emerges from a general dynamic system with diverse microscopic interactions.
@@
As the first step towards understanding how relativistic nonequilibrium systems reach thermal equilibrium, the properties of retarded correlation functions have recently garnered extensive research interests. Two-point retarded correlation functions are pivotal, as they provide a wealth of insights into the transport characteristics of multi-particle systems, particularly how an equilibrium system responds to off-equilibrium disturbances within the linear regime. Moreover, non-analytical structures—such as poles or discontinuities in Fourier space—are crucial for determining the system’s evolutionary patterns. Poles are indicative of collective excitations that evolve towards equilibrium, corresponding to the persistent hydrodynamic modes, whereas the presence of cuts or non-hydrodynamic modes is intrinsically linked to the emergence of hydrodynamic phenomena and the applicability of hydrodynamics. Research has shown that the correlators contain only poles at infinite t’Hooft coupling in large $N$ thermal $\mathcal{N}=4$ Super Yang-Mills (SYM) theory. In the holographic description, the spectra of correlation functions correspond to the ring-down spectra of dual linearly perturbed black holes: the quasinormal modes [5, 6]. However, in Ref. [5], the authors argue that the branch-cut structure emerges in the regime of weak, but finite t’Hooft coupling, indicating a possible transition behavior controlled by the t’Hooft coupling. Motivated by this illuminating finding, Paul Romatschke initiated research on thermal correlators of large $N$ gauge theories in effective kinetic theory [7](see also [8]), reporting the onset of transition behavior for hydrodynamic poles. In subsequent related studies, including analytical estimates [9], qualitative models [10, 11], and numerical calculations [12, 13], the dominant non-analytical structure is found to be the branch-cut rather than poles, thus posing the ’poles or cuts?’ dilemma. Specifically, this dilemma initially refers to the mathematical essence of non-hydrodynamic excitations. In this work, we also use it to the denote the comparison of the lifetime of hydrodynamic/non-hydrodynamic excitations. This is one of the topics we aim to explore in this manuscript, corresponding to Q2 as introduced in the abstract.
@@
However, the Boltzmann equation still contains an intractable collision integral, rendering its analytical or numerical solution challenging. Even with the simplest interactions, such as the hard-sphere potential, the linearized Boltzmann equation’s collision operator retains a complex structure, making its analysis extremely difficult. Recently, the eigenspectrum of the linearized Boltzmann collision operator in massless scalar $\phi^{4}$ theory was analytically determined [18]. However, extending this to other realistic interactions remains a significant challenge [^1] The first analytical eigenspectrum was obtained by C.S.Wang Chang and U.E.Uhlenbeck in the context of monatomic gases in the non-relativistic case (see chapter IV of [19]).. In 1974, Anderson and Witting (AW) proposed an approximation of the relativistic Boltzmann equation using a simplified collision operator [20], effectively extending the BGK (Bhatnagar, Gross, and Krook) model to the relativistic domain [21]. The AW model, also known as the relaxation time approximation, omits much of the dynamic information in the full collision operator. Firstly, it disregards nonlinearity. Secondly, its validity hinges on a clear separation between the eigenvalue representing the slowest relaxation and the others. Despite these limitations, the relaxation time approximation addresses practical challenges, enabling analytic and semi-analytic solutions for the simplified Boltzmann equation. This approach paves the way for an insightful and instructive analysis. Furthermore, the AW RTA model has been extended to a more general form with nontrivial energy dependence. However, the justification of this extended RTA model within kinetic theory awaits rigorous validation, which is another topic of interest, referred to as Q1 in the abstract.
[…]
@@ II Linearized Boltzmann equation
$-\mathcal{L}_{0}[\chi]$ $\equiv E_{p}^{-1}\int{\rm dP^{\prime}}{\rm dP_{1}}{\rm dP_{2}}f_{0}(x,p^{\prime})W_{p,p^{\prime}\to p_{1},p_{2}}$
$\big{(}\chi(x,p_{1})+\chi(x,p_{2})-\chi(x,p)-\chi(x,p^{\prime})\big{)},$ (4)
@@ III Revisiting the relaxation time approximation
where $\mathcal{L}_{\alpha}\equiv E_{p}^{\alpha}\mathcal{L}_{0}$ represents the redefined linearized collision operator with $\mathcal{L}_{1}$ and $\mathcal{L}_{0}$ being specific instances. It is easily proved that $\mathcal{L}_{\alpha}$ inherits the positive semidefinite and self-adjoint properties from $\mathcal{L}_{0}$. This can be achieved by redefining the weight function within the inner product definition as $f_{0}(x,p)E_{p}^{1-\alpha}$ in Eq.(II). In the following subsections III.2, III.2 and III.3, we illustrate that naively truncating $\mathcal{L}_{\alpha}$ leads to a form resembling the energy-dependent RTA; however the resulting RTA-like model is not well-defined.
@@ III.1 Anderson and Witting model
For comparative purposes, we also present the expansion form of $\chi^{\prime}=\sum_{n}c^{\prime}_{n}(t)\psi_{n}(p)$. Additionally, the counter term $-\sum_{n=1}^{5}c^{\prime}_{n}(0)\psi_{n}(p)e^{-t/\tau_{R}}$ is introduced to account for the absence of collision invariance in the traditional RTA, as further discussed in Sec.V. Note $C(p),C^{\prime}(p)$ are integration constants. In the above equations, the eigenvalues are sequenced in ascending order when increasing $n$. As clearly seen in Eq.(14), the collision invariants, $\psi_{n}(n=1,\cdots 5)$ with zero eigenvalues $\gamma_{n}=0$, contribute to the deviation function $\chi$ but remain unchanged with time. In the solution of Eq.(14), the second term, reflecting the contribution from collision invariants, can also be regarded as an integration constant. By comparing Eqs.(14) and (15), we can deduce that the traditional RTA simplifies the model by condensing the entire nonzero eigenvalue spectrum into a single representative eigenvalue, ${\gamma_{n}}\rightarrow\frac{1}{\tau_{R}}$. Given the exponentially decaying form, $1/\tau_{R}$ should be identified as the smallest eigenvalue, where the mode $\psi_{6}$ persists until the late stage of evolution. In other words, other modes with a larger eigenvalue, which should have been absent from an earlier time, extend their lifetime to the final stage in this approximation.
However, the RTA model is justified only when the eigenvalue sequence ${\gamma_{n}}$ is distinctly separated from the origin. If this is not the case, such as when the eigenvalue spectrum extends continuously from the origin to other nonzero points, the relaxation time cannot be defined as the inverse of the smallest eigenvalue, as it might diverge, leading to the elimination of collisional effects. Thus, the justification of the RTA heavily depends on the eigenspectrum structure of the linearized Boltzmann collision operator, particularly the characteristics of the eigenspectrum near the origin. As shown in III.4 of this section, the traditional RTA is permissible only in scenarios of ’hard-interaction’ collisions, where the eigenspectrum of the linearized collision operator features a continuous range from $\nu_{0}\,(\nu_{0}>0)$ to $\infty$, along with possibly some discrete points. In such scenarios, a gap exist between the origin and other nonzero eigenvalues, providing a basis for constructing the RTA.
@@ III.2 Energy-dependent relaxation time approximation ($\alpha>0$)
As noted in the convenient parameterization Eq.(8), the relaxation time, the only parameter reflecting the microscopic dynamic details within the RTA model, is allowed to exhibit energy dependence. It has long been believed that incorporating various energy dependence can reveal characteristic features of bottom-up thermalization and uncover hidden aspects of the full kinetic description [10]. Furthermore, the linearized kinetic equation within the energy-dependent relaxation time approximation has been successfully applied to jet physics [28, 29], hydrodynamic simulations [25, 26], and Weyl semimetals [30]. Does this imply that the model or approximation is well-justified? Caution should be taken: an approximated model may capture some aspects of the underlying physics relevant to specific issues but may fail to do so for others. Nevertheless, elucidating how an approximation is derived from a complete theory is always highly significant. In this subsection, we show mathematically how this model can be developed from the linearized Boltzmann in a similar manner to what we have presented in the former subsection, and how an inconsistence appears hindering the sound justification of the model.
@@
$\chi^{\prime}(t,p)$ $=\sum_{n}d_{n}^{\prime}(0)\Psi_{n}(p)e^{-t/(\beta E_{p}^{\alpha}t_{R})}-\sum_{n=1}^{5}d_{n}^{\prime}(0)\Psi_{n}(p)e^{-t/(\beta E_{p}^{\alpha}t_{R})}+D^{\prime}(p)=\sum_{n>5}d_{n}^{\prime}(0)\Psi_{n}(p)e^{-t/(\beta E_{p}^{\alpha}t_{R})}+D^{\prime}(p),$ (19)
where $\{\gamma^{\prime}_{n},\Psi_{n}\}$ represents the eigensystem of $\mathcal{L}_{\alpha>0}=E^{\alpha}_{p}\mathcal{L}_{0}$, and $D(p),D^{\prime}(p)$ are integration constants. Given that $E_{p}^{\alpha}\geq 0$, it still holds that the eigen modes with a larger eigenvalue fade away more quickly. Depending on whether the eigenspectrum continuously extends to the origin, we decide to retain the smallest eigenvalue but zero or not. It seems that the story has no changes at all compared to what have been done in the previous subsection.
@@
Note as an aside, $\alpha$ cannot assume arbitrary values for massless particles, as this could lead to an ill-defined inner product. Since the collision invariant $1$ must be included in the Hilbert space, $\alpha$ should not exceed $4$ to avoid infrared divergence in the integral
[…]
This statement is made under the assumption that the discussed functions, such as the eigenfunctions of $\mathcal{L}_{\alpha}$, behave regularly in the finite $p$ region, excluding singular behaviors like $\sim\frac{1}{p}$.
@@ III.3 Energy-dependent relaxation time approximation ($\alpha<0$)
The necessity of the third condition can be relaxed if we deviate from the widely used parameterization in Eq.(8). Additionally, the second condition may also be loosed by employing alternative parameterizations. In this analysis, we focus our attention on the parameterization specified in Eq.(8).
@@
We ’naively’ truncate $\mathcal{L}_{\alpha<0}$ to the smallest nonzero eigenvalue $\gamma^{\prime}_{6}$ (or the nonzero infimum if the spectrum is continuous) to reproduce Eq.(8). Thus, the relaxation timescale for the perturbation decay is precisely $\frac{E_{p}^{\alpha}}{\gamma^{\prime}_{6}}$ bounded from above by $\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$. This indicates that all modes attenuate at a finite rate with the slowest rate interpreted as the inverse relaxation time, $1/\tau_{R}=\frac{\gamma^{\prime}_{6}}{m^{\alpha}}$, in contrast to what we encounter in III.2. But if the particles are massless, then it follows that $\frac{E_{p}^{\alpha}}{\gamma^{\prime}_{6}}$ is unbounded, and the well-defined RTA is also lacking for the similar reason shown in III.2.
However, we confirm that $\tau_{R}=\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$ equates to the scenario where $\alpha=0$, i.e., the AW RTA, and does not introduce a new class of RTAs. Suppose an observer perturbs the system in equilibrium with a small-amplitude spatially homogeneous disturbance at $t=0$, then the observer is observing when the disturbance dissipates to vanishing. For the system described by Eq.(21), this relaxation scale is dictated by the mode with the longest lifetime, and reads $\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$. Hence at $t=\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$, the observer announces that the system returns to its equilibrium state. If the same experiment is conducted for the system Eq.(12), the observer should get the same result, as Eq.(21) is an equivalent transformation of Eq.(12). For future convenience in expression, this argument will be also referred to as the “observer argument”. In other words, if condition 1 is satisfied, $\mathcal{L}_{0}$ also possesses nonzero eigenvalues separated from the origin, and these two relaxation scales should be identical, $\frac{1}{\gamma_{6}}\simeq\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$. Several comments on this relation are followed in order:
  • We identify a gapped scale $\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$, which bounds the relaxation timescales for perturbation decay. This means a RTA description emerges, with the relaxation time as given in Eq.(8) but with $E_{p}\rightarrow m$, which reminds me of an inverse but interesting replacement of $m\rightarrow E_{p}$ in the initial construction for AW RTA [20]. But ruled by the above physical arguments, this deformed energy-dependent RTA (precisely mass-dependent RTA) is still a redundant description. In this sense, we conclude that there is no energy-dependent RTA for $\mathcal{L}_{\alpha<0}$.
  • The interesting relation $\frac{1}{\gamma_{6}}\simeq\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$ emerges purely from physical considerations without any mathematical analysis. The premise that an observer should detect identical relaxation behaviors could be extended to other scenarios. For example, if $\mathcal{L}_{0}$ has only a discrete eigenspectrum or a gapped continuous eigenspectrum in the vicinity of origin so that the RTA becomes feasible, how would the eigenspectrum of $\mathcal{L}_{\alpha>0}$ be ? We shall come to this question later.

### v3

# Relaxation time approximation revisited and non-analytical structure in retarded correlators

@@ Abstract
In this paper, we give a rigorous mathematical justification for the relaxation time approximation (RTA) model. We find that only the RTA with an energy-independent relaxation time can be justified in the case of hard interactions. Accordingly, we propose an alternative approach to restore the collision invariance lacking in traditional RTA. Besides, we provide a general statement on the non-analytical structures in the retarded correlators within the kinetic description. For hard interactions, hydrodynamic poles are the long-lived modes. Whereas for soft interactions, commonly encountered in relativistic kinetic theory, the gapless eigenvalue spectrum of linearized collision operator leads to gapless branch-cuts. We note that particle mass and inhomogeneous perturbations would complicate the above-mentioned non-analytical structures.
@@ I Introduction
Over the past several decades, relativistic hydrodynamics has notably contributed to characterizing the dynamic evolution of Quark-Gluon Plasma (QGP), as observed in heavy-ion collision experiments at the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) [1]. Within the framework of phenomenological modeling for these collisions, the initialization time of hydrodynamic simulation is usually set to be less than $1$ fm/c to match experimental observations. This suggests that the system may rapidly equilibrate. Moreover, the applicability of hydrodynamics has been noted to extend into less expected domains, including small collision systems like nucleus-nucleon and proton-proton collisions [2, 3], which implies that these small collision systems could possibly exhibit fluid-like behaviors. What is the dynamic mechanism that triggers rapid equilibration? Why does relativistic hydrodynamics also work well for small systems? Regarding the first question, it has been proposed that hydrodynamics can be viewed as an attractor governing the late-time behavior of systems as they approach equilibrium [4]. At the early stage of evolution, the system tends to flow towards the hydrodynamic attractor, even when far from local equilibrium, which may account for rapid hydrodynamization. Thus, hydrodynamization may have a broader application range than thermalization, contrasting with the conventional view that hydrodynamics is a truncated gradient expansion near local equilibrium. Despite significant advancements, it remains an open question how hydrodynamization emerges from a general dynamic system with diverse microscopic interactions.
@@
As the first step towards understanding how relativistic nonequilibrium systems reach thermal equilibrium, the properties of retarded correlation functions have recently garnered extensive research interests. Two-point retarded correlation functions are pivotal, as they provide a wealth of insights into the transport characteristics of multi-particle systems, particularly how an equilibrium system responds to off-equilibrium disturbances within the linear regime. Moreover, non-analytical structures—such as poles or discontinuities in Fourier space—are crucial for determining the system’s evolutionary patterns. Poles are indicative of collective excitations that evolve towards equilibrium, corresponding to the persistent hydrodynamic modes, whereas the presence of non-hydrodynamic modes, which could be cuts or poles, is intrinsically linked to the emergence of hydrodynamic phenomena and the applicability of hydrodynamics. Research has shown that the correlators contain only poles at infinite ’t Hooft coupling in large $N$ thermal $\mathcal{N}=4$ Super Yang-Mills (SYM) theory. In the holographic description, the spectra of correlation functions correspond to the ring-down spectra of dual linearly perturbed black holes: the quasinormal modes [5, 6]. However, in Ref. [5], the authors argue that the branch-cut structure emerges in the regime of weak, but finite ’t Hooft coupling, indicating a possible transition behavior controlled by the ’t Hooft coupling. Motivated by this illuminating finding, Paul Romatschke calculated thermal correlators of large $N$ gauge theories in effective kinetic theory [7] (see also [8, 9]), reporting the onset of transition behavior for hydrodynamic poles. In other related studies, including analytical estimates [10], qualitative models [11, 12], and numerical calculations [13, 14], the dominant non-analytical structure is found to be the branch-cut rather than poles, thus posing the ’poles or cuts?’ dilemma. Specifically, this dilemma initially refers to the mathematical essence of non-hydrodynamic excitations. In this work, we also use it to denote the comparison of the lifetime of hydrodynamic/non-hydrodynamic excitations. Also, it is worth noting that in a recent study [15] on the transient dynamics of quasinormal modes within a holographic framework, the authors demonstrated the existence of arbitrarily long-lived excitations arising from summation of short-lived quasinormal modes. These correspond to initial conditions in which a packet of energy is localized near the future horizon, a feature that bears a resemblance to the numerical results presented in [13, 14]. This may suggest a deeper and more direct connection between the linear response behaviors described by weakly-coupled kinetic theory and strongly-coupled holographic theories.
@@
However, the Boltzmann equation still contains an intractable collision integral, rendering its analytical or numerical solution challenging. Even with the simplest interactions, such as the hard-sphere potential, the linearized Boltzmann equation’s collision operator retains a complex structure, making its analysis extremely difficult, see [20] for related discussions on the linearized collision operator spectrum for hard-sphere interaction. Recently, the eigenspectrum of the linearized Boltzmann collision operator in massless scalar $\phi^{4}$ theory was analytically determined [21]. However, extending this to other realistic interactions remains a significant challenge [^1] The first analytical eigenspectrum was obtained by C.S.Wang Chang and U.E.Uhlenbeck in the context of monatomic gases in the non-relativistic case (see chapter IV of [22]).. Marle [23, 24], and Anderson and Witting (AW) [25] proposed an approximation of the relativistic Boltzmann equation using a simplified collision operator, effectively extending the BGK (Bhatnagar, Gross, and Krook) model to the relativistic domain [26]. In the case of Marle, the proportionality factor is given by $m/\tau_{R}$, with $m$ the mass of the particles and $\tau_{R}$ the relaxation time, whereas for AW model, it is $u\cdot p/\tau_{R}$. We focus on the latter case in this work. The AW model, also known as the relaxation time approximation, omits much of the dynamic information in the full collision operator. Firstly, it disregards nonlinearity. Secondly, its validity hinges on a clear separation between the eigenvalue representing the slowest relaxation and the others. Despite these limitations, the relaxation time approximation addresses practical challenges, enabling analytic and semi-analytic solutions for the simplified Boltzmann equation. This approach paves the way for an insightful and instructive analysis. Furthermore, the AW RTA model has been extended to a more general form with nontrivial energy dependence. However, the justification of this extended RTA model within kinetic theory awaits rigorous validation, which is another topic of interest.
The paper is organized as follows. In Sec. II, we briefly review the basic aspects of the linearized Boltzmann equation. In Sec. III, we revisit the formulation of relaxation time approximation within the linearized Boltzmann equation. In this section, we provide a justification for the RTA, based on general mathematical considerations regarding the eigenspectrum structure of the linearized collision operator. It turns out that the RTA is well-justified exclusively in scenarios involving relativistic hard interactions. Sec. IV serves as an application of the findings from Sec. III to address the non-analytical structures contained in the retarded correlation functions. In Sec. V, we introduce a novel relaxation time approximation by truncating the full linearized operator, anticipating a broader range of applicability. Summary and outlook are given in Sec. VI.
Natural units $k_{B}=c=\hbar=1$ are employed. The metric tensor is given by $g^{\mu\nu}=\operatorname{diag}(1,-1,-1,-1)$ , while $\Delta^{\mu\nu}\equiv g^{\mu\nu}-u^{\mu}u^{\nu}$ is the projection tensor orthogonal to the four-vector fluid velocity $u^{\mu}$. The abbreviation dP stands for $\int dP\equiv\frac{2}{(2\pi)^{3}}\int d^{4}p\,\theta(p^{0})\delta(p^{2}-m^{2})$.
@@ II Linearized Boltzmann equation
$-\mathcal{L}_{0}[\chi]$ $\equiv E_{p}^{-1}\int{\rm dP^{\prime}}{\rm dP_{1}}{\rm dP_{2}}f_{0}(x,p^{\prime})W_{p,p^{\prime}\to p_{1},p_{2}}\big(\chi(x,p_{1})+\chi(x,p_{2})-\chi(x,p)-\chi(x,p^{\prime})\big),$ (4)
@@ III Revisiting the relaxation time approximation
where $\mathcal{L}_{\alpha}\equiv E_{p}^{\alpha}\mathcal{L}_{0}$ represents the redefined linearized collision operator with $\mathcal{L}_{1}$ and $\mathcal{L}_{0}$ being specific instances. It is easily proved that $\mathcal{L}_{\alpha}$ inherits the positive semidefinite and self-adjoint properties from $\mathcal{L}_{0}$. This can be achieved by redefining the weight function within the inner product definition as $f_{0}(x,p)E_{p}^{1-\alpha}$ in Eq.(II), which means the square-integrable function space should be also altered accordingly. It should be noted that the corresponding integrals within the definitions should converge, which excludes very large values for $\alpha$. In the following subsections III.2 and III.3, we illustrate that naively truncating $\mathcal{L}_{\alpha}$ leads to a form resembling the energy-dependent RTA; however the resulting RTA-like model is not well-defined.
@@ III.1 Anderson and Witting model
For comparative purposes, we also present the expansion form of $\chi^{\prime}=\sum_{n}c^{\prime}_{n}(t)\psi_{n}(p)$. Additionally, the counter term $\sum_{n=1}^{5}c^{\prime}_{n}(0)\psi_{n}(p)e^{-t/\tau_{R}}$ is introduced to account for the absence of collision invariance in the traditional RTA, as further discussed in Sec.V. Note $C(p),C^{\prime}(p)$ are integration constants. In the above equations, the eigenvalues are sequenced in ascending order when increasing $n$. As clearly seen in Eq.(14), the collision invariants, $\psi_{n}(n=1,\cdots 5)$ with zero eigenvalues $\gamma_{n}=0$, contribute to the deviation function $\chi$ but remain unchanged with time. In the solution of Eq.(14), the second term, reflecting the contribution from collision invariants, can also be regarded as an integration constant. By comparing Eqs.(14) and (15), we can deduce that the traditional RTA simplifies the model by condensing the entire nonzero eigenvalue spectrum into a single representative eigenvalue, ${\gamma_{n}}\rightarrow\frac{1}{\tau_{R}}$. Given the exponentially decaying form, $1/\tau_{R}$ should be identified as the smallest eigenvalue, where the mode $\psi_{6}$ persists until the late stage of evolution. In other words, other modes with a larger eigenvalue, which should have been absent from an earlier time, extend their lifetime to the final stage in this approximation.
However, the RTA model is justified only when the eigenvalue sequence $\gamma_{n}$} is distinctly separated from the origin. If this is not the case, such as when the eigenvalue spectrum extends continuously from the origin to other nonzero points, the relaxation time cannot be defined as the inverse of the smallest eigenvalue, as it might diverge, leading to the elimination of collisional effects. Thus, the justification of the RTA heavily depends on the eigenspectrum structure of the linearized Boltzmann collision operator, particularly the characteristics of the eigenspectrum near the origin. As shown in III.4 of this section, the traditional RTA is permissible only in scenarios of ’hard-interaction’ collisions, where the eigenspectrum of the linearized collision operator features a continuous range from $\nu_{0}\,(\nu_{0}>0)$ to $\infty$, along with possibly some discrete points. In such scenarios, a gap exists between the origin and other nonzero eigenvalues, providing a basis for constructing the RTA.
@@ III.2 Energy-dependent relaxation time approximation ($\alpha>0$)
As noted in the convenient parameterization Eq.(8), the relaxation time, the only parameter reflecting the microscopic dynamic details within the RTA model, is allowed to exhibit energy dependence. It has long been believed that incorporating various energy dependence can reveal characteristic features of bottom-up thermalization and uncover hidden aspects of the full kinetic description [11]. Furthermore, the linearized kinetic equation within the energy-dependent relaxation time approximation has been successfully applied to jet physics [33, 34], hydrodynamic simulations [30, 31], and Weyl semimetals [35]. Does this imply that the model or approximation is well-justified? One caution should be taken: an approximated model may capture some aspects of the underlying physics relevant to specific issues but may fail to do so for others. Nevertheless, elucidating how an approximation is derived from a complete theory is always highly significant. In this subsection, we show mathematically how this model can be similarly derived from the linearized Boltzmann equation, and how an inconsistency appears hindering the sound justification of the model.
@@
$\chi^{\prime}(t,p)$ $=\sum_{n}d_{n}^{\prime}(0)\Psi_{n}(p)e^{-t/(\beta E_{p}^{\alpha}t_{R})}-\sum_{n=1}^{5}d_{n}^{\prime}(0)\Psi_{n}(p)e^{-t/(\beta E_{p}^{\alpha}t_{R})}+D^{\prime}(p)$
$=\sum_{n>5}d_{n}^{\prime}(0)\Psi_{n}(p)e^{-t/(\beta E_{p}^{\alpha}t_{R})}+D^{\prime}(p),$ (19)
where $\{\gamma^{\prime}_{n},\Psi_{n}\}$ represents the eigensystem of $\mathcal{L}_{\alpha>0}=E^{\alpha}_{p}\mathcal{L}_{0}$, and $D(p),D^{\prime}(p)$ are integration constants. Given that $E_{p}^{\alpha}\geq 0$, it still holds that the eigen modes with a larger eigenvalue fade away more quickly. Depending on whether the eigenspectrum continuously extends to the origin, we decide to retain the smallest eigenvalue but zero or not. It seems that the story remains no changes at all compared to what has been done in the previous subsection.
@@ III.3 Energy-dependent relaxation time approximation ($\alpha<0$)
The necessity of the third condition can be relaxed if we deviate from the widely used parameterization in Eq.(8). Additionally, the second condition may also be loosened by employing alternative parameterizations. In this analysis, we focus our attention on the parameterization specified in Eq.(8).
@@
We “naively” truncate $\mathcal{L}_{\alpha<0}$ to the smallest nonzero eigenvalue $\gamma^{\prime}_{6}$ (or the nonzero infimum if the spectrum is continuous) to reproduce Eq.(8). Thus, the relaxation timescale for the perturbation decay is precisely $\frac{E_{p}^{\alpha}}{\gamma^{\prime}_{6}}$ bounded from above by $\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$. This indicates that all modes attenuate at a finite rate with the slowest rate interpreted as the inverse relaxation time, $1/\tau_{R}=\frac{\gamma^{\prime}_{6}}{m^{\alpha}}$, in contrast to what we encounter in III.2. But if the particles are massless, then it follows that $\frac{E_{p}^{\alpha}}{\gamma^{\prime}_{6}}$ is unbounded, and the well-defined RTA is also lacking for the similar reason shown in III.2.
However, we confirm that $\tau_{R}=\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$ equates to the scenario where $\alpha=0$, i.e., the AW RTA, and does not introduce a new class of RTAs. This verification is ruled by the following physical argument. Suppose an observer perturbs the system in equilibrium with a small-amplitude spatially homogeneous disturbance at $t=0$, then the observer records when the disturbance dissipates to none. It is reasonable to require that Eqs.(3), (9), and their corresponding RTAs (if any) give the same relaxation timescale.
Assuming that the eigenvalue spectrum of $\mathcal{L}_{0}$ is gapped, the relaxation scale is given by $\frac{1}{\gamma_{6}}$ according to Eq.(3). If we try to describe the system using Eq.(9) with $\alpha<0$, the relaxation scale is then dictated by the mode with the longest lifetime, and reads $\tau_{R}(p\rightarrow 0)=\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$. Hence at $t=\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$, the observer announces that the system returns to its equilibrium state. Given that the same experiment is conducted for the same system, the observer should give identical observation results, as Eq.(9) is an equivalent transformation of Eq.(3), i.e., $\frac{1}{\gamma_{6}}\simeq\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$. For the $\alpha<0$ case, even if there is an upper bound for relaxation scales and one can manufacture an energy-dependent (or precisely mass-dependent) RTA, it is still a redundant description compared to AW RTA. In summary, an energy-dependent RTA should not be regarded as a well-defined truncation to its UV complete theory, i.e., the linearized Boltzmann equation, irrespective of the interaction details. It is suggestive to keep only the description provided by $\mathcal{L}_{0}$, because it is simplest without energy-dependent factor, and the mathematical analysis specifically for $\mathcal{L}_{0}$ is readily available [36, 37].
The observer argument above can provide an equivalent justification for reconciling the seemingly contradictory aspects between two detailed calculations for massless scalar $\phi^{4}$ theory with leading order interaction [14, 21]. Denicol and Noronha analytically calculate the eigenspectrum of $\mathcal{L}_{1}$ and obtain a series of discrete eigenvalues [21]. By utilizing the observer argument again, we can conclude that the eigenspectrum of $\mathcal{L}_{0}$ must extend continuously towards the origin, and thus gapless. This conclusion is consistent with the mathematical aspect: one can show that for leading order scalar $\phi^{4}$ theory, $\sigma(g,\Theta)\sim\frac{1}{s}=\frac{1}{g^{2}+4m^{2}}<\frac{2}{g^{2}}$, so the interaction falls within the soft interaction category according to the criteria given in [36, 37]. Furthermore, the gapless continuous eigenspectrum has also been confirmed in recent calculations [14, 27, 38].